# HybridFlow: Adaptive Task Scheduling for Fast and Token-Efficient LLM Inference in Edge-Cloud Collaboration

### v2

@@ Abstract
Large language models (LLMs) exhibit impressive reasoning and problem-solving abilities, yet their substantial inference latency and token consumption pose major challenges for real-time deployment on resource-limited edge devices. Recent efforts toward edge-cloud collaboration have attempted to mitigate this issue, but most existing methods adopt coarse-grained task allocation strategies-assigning entire queries either to the edge or the cloud. Such rigid partitioning fails to exploit fine-grained reasoning parallelism and often leads to redundant computation and inefficient resource utilization. To this end, we propose HybridFlow, a resource-adaptive inference framework that enables fast and token-efficient collaborative reasoning between edge and cloud LLMs. HybridFlow operates in two stages: (1) task decomposition and parallel execution, which dynamically splits a complex query into interdependent subtasks that can execute as soon as their dependencies are resolved; and (2) resource-aware subtask routing, where a learned router adaptively assigns each subtask to the edge or cloud model according to predicted utility gains and real-time budget states. Comprehensive evaluations on GPQA, MMLU-Pro, AIME, and LiveBench-Reasoning demonstrate that HybridFlow effectively reduces end-to-end inference time and overall token usage while maintaining competitive accuracy. Our code is available at: https://github.com/Jiangwen-Dong/Hybridflow-1
@@ 2 Related Work
A growing line of work improves LLM reasoning efficiency by restructuring inference into smaller units. CoT-style prompting (Wei et al., 2022) enhances accuracy but often increases token and latency costs due to long intermediate traces. More recent approaches introduce structured decomposition and partial parallelism, such as SoT (Ning et al., 2024) and PASTA (Jin et al., 2025), which elicit intermediate sub-questions or steps that may be executed concurrently. Beyond linear decompositions, ToT (Yao et al., 2023) organizes reasoning as a tree and performs deliberate search over candidate thoughts, while GoT (Besta et al., 2024) and S-DAG (Dong et al., 2025) generalizes this idea to graph-structured reasoning with reusable intermediate nodes and flexible dependency patterns.
Budgeted Edge–Cloud Routing with Online Adaptation.
Edge–cloud collaboration commonly relies on routing mechanisms, where a small model handles easy cases and offloads harder ones to a stronger model, or where a task is decomposed and allocated across heterogeneous models (Wang et al., 2025; Li et al., 2025; Ding et al., 2024; Shao et al., 2025). Many such systems operate at coarse granularity (e.g., query-level or stage-level) and do not explicitly optimize allocations under tight, multi-resource constraints such as latency and API budgets. In parallel, cost-aware cascade and routing frameworks aim to meet a target quality under budget by selectively invoking stronger models; for example, LLM cascades based on mixture-of-thought style representations (Yue et al., 2023) and FrugalGPT (Chen et al., 2023) emphasize budget–quality trade-offs via staged model selection and routing. More closely to fine-grained offloading, SplitReason (Akhauri et al., 2025) learns to offload parts of reasoning to a stronger model, highlighting the benefit of step-level delegation. However, many learned routers assume stationary conditions or fixed thresholds, which can be brittle under shifts in workload, network latency, or pricing.
Algorithm 1 HybridFlow: Adaptive Task Planning and Resource-Aware Routing
Input: User query $Q$; global API budget $K_{\max}^{\mathrm{global}}$; global latency budget $L_{\max}^{\mathrm{global}}$
@@
Stage 1: Task Decomposition
$T\leftarrow\mathrm{Decompose}(Q;M_{P})$
Construct DAG $G=(T,E)$ and initialize queue $\mathcal{Q}$ with all dependency-free subtasks.
Stage 2: Resource-Aware Routing and Execution
Initialize resource usage $K_{\mathrm{used}}=0$, $L_{\mathrm{used}}=0$.
while $\mathcal{Q}$ not empty do
   Pop subtask $x$ from $\mathcal{Q}$. Obtain subtask embedding: $z(x)=\texttt{embedding}(x)$ Predict utility score $\hat{u}(x)=f_{\theta}(z(x))$. Compute adaptive threshold $\tau_{t}$ from current resource usage. if $\hat{u}(x)>\tau_{t}$ then
    $R_{x}\leftarrow M_{\mathrm{cloud}}(x)$; ;
    // Cloud executionUpdate $K_{\mathrm{used}}$ and $L_{\mathrm{used}}$.
@@
    $R_{x}\leftarrow M_{\mathrm{edge}}(x)$; ;
    // Edge executionUpdate $L_{\mathrm{used}}$.
  Append $R_{x}$ to global context and push newly unblocked subtasks into $\mathcal{Q}$.
Stage 3: Final AggregationCombine all sub-results in topological order to form $R$. return $R$
@@ 3 HybridFlow Framework
HybridFlow enables efficient edge–cloud collaborative inference by combining dependency-aware task planning with cost-aware routing. As illustrated in Fig. 1, the system has two tightly connected components. First, an edge-side planner decomposes a complex query into a set of interdependent subtasks and constructs a task-level DAG, enabling each subtask to execute as soon as its prerequisites are resolved. This exposes fine-grained parallelism and reduces wall-clock latency compared with strictly sequential reasoning. Second, a resource-aware router assigns each subtask to either the edge model $M_{\mathrm{edge}}$ or the cloud LLM $M_{\mathrm{cloud}}$ based on its predicted utility and the system’s real-time budget status. By coupling decomposition with budget-constrained routing, HybridFlow explicitly balances accuracy and efficiency.
@@ 4.3 Ablation Study
We evaluate the effectiveness of our adaptive router by comparing it against several deterministic and naive allocation strategies. As shown in Table 3, the results reveal a clear trade-off between accuracy and resource cost. Using only the edge model (Edge) yields the lowest accuracy (25.54%) but incurs zero API cost, while delegating all subtasks to the cloud model (Cloud) achieves the highest accuracy (57.28%) at the expense of higher latency and API cost ($C_{\text{API}}{=}0.0185$). A random assignment baseline (Random) moderately improves accuracy (46.00%) by occasionally leveraging the cloud model but still wastes resources and fails to adapt to budget constraints. In contrast, our HybridFlow achieves 53.33% accuracy with a comparable API cost ($C_{\text{API}}{=}0.0075$), demonstrating its ability to intelligently allocate subtasks according to their benefit–cost ratio. This shows that the learned router effectively captures the trade-off between reasoning quality and resource expenditure, achieving near-optimal performance under tight cost budgets.
@@
To further investigate the router’s allocation behavior within a single reasoning process, we record the number of subtasks executed on the edge and in the cloud at each subtask position. As illustrated in Figure 3, the offload pattern is highly structured rather than uniform across positions. For the first few subtasks, when the adaptive threshold is low and the resource budget is still largely unused, the router frequently offloads to the cloud, reflecting a preference to invest cloud capacity in early, semantically complex steps that shape the overall solution. As the reasoning progresses, the threshold steadily increases, and we observe a clear shift: later subtasks are predominantly executed on the edge, with very few cloud calls once the threshold saturates. At the same time, the absolute number of subtasks also decreases at deeper positions, indicating that most plans resolve the main reasoning bottlenecks early. Together, these trends confirm that HybridFlow does not offload subtasks indiscriminately, but instead allocates cloud resources to earlier, high-impact steps and relies on the edge model for downstream, lower-utility subtasks as the budget tightens.
@@
To further analyze the router’s behavior, we vary the fixed base threshold $\tau_{0}$ in Eq. 11 (without dynamic resource adaptation) and measure its effect on accuracy, cost, and efficiency. As shown in Figure 4 and Table 5, increasing $\tau_{0}$ makes the router more conservative, progressively reducing both offload rate and normalized cost (from 100% and 0.92 at $\tau_{0}{=}0$ to nearly 0 at $\tau_{0}{=}1$), while accuracy decreases smoothly from 57.28% to 25.54%. This monotonic trend indicates that the router’s predicted utility scores $\hat{u}_{i}\in[0,1]$ are well aligned with the normalized cost scale $c_{i}\in[0,1]$, ensuring coherent decision boundaries across thresholds.
Interestingly, the utility score (purple curve) peaks around $\tau_{0}{=}0.6$, corresponding to the most balanced regime where accuracy ($\approx$52.61%) is maintained at roughly half the total cost (c=0.45c{=}0.45). As also shown in Table 3, HybridFlow achieves an even higher utility (0.6294), indicating that moderate thresholds yield the most efficient trade-off. Overall, the fixed-threshold analysis confirms that the router normalized benefit–cost predictions are coherent and interpretable, producing stable and efficient allocations while avoiding the extremes of excessive offloading or underutilization.
@@ Appendix A Notations
$\tau_{t}$ Adaptive routing threshold at time $t$ [Eq. (LABEL:equ:threshold)].
@@ B.3 Primal–Dual Dynamics and Adaptive Thresholding
HybridFlow’s adaptive threshold in Eq. (LABEL:equ:threshold) of the main text,

### v3

# HybridFlow: Resource-Adaptive Subtask Routing for Efficient Edge-Cloud LLM Inference

@@ Abstract
Edge-cloud collaborative inference is becoming a practical necessity for LLM-powered edge devices: on-device models often cannot afford the required reasoning capability, while cloud-only inference could be prohibitively costly and slow under strict latency and token/API budgets. However, existing edge-cloud collaboration methods often route per query or fixed steps simply based-on the estimated difficulty. Such coarse and static heuristics overlook subtask dependencies, missing opportunities for parallel execution and budget-adaptive routing. To this end, we propose HybridFlow, a resource-adaptive edge-cloud inference framework that (i) builds a dependency-aware DAG for each query and executes newly unlocked subtasks in parallel, reducing end-to-end latency; (ii) routes each subtask online to the edge or cloud via a learned benefit–cost utility model that dynamically trades accuracy gains against token/API and latency budgets, thereby reducing unnecessary cloud usage while preserving reasoning quality. Across GPQA, MMLU-Pro, AIME24, and LiveBench-Reasoning, HybridFlow improves the cost-accuracy trade-off, reducing latency and cloud API usage while maintaining competitive accuracy against strong structured reasoning baselines. Code: HybridFlow
@@ 2 Related Work
A growing body of research aims to improve LLM reasoning efficiency by decomposing inference into smaller, more manageable units. CoT prompting (Wei et al., 2022) enhances reasoning accuracy but often increases token and latency costs due to long intermediate chains. More recent works have introduced structured decomposition and partial parallelism to address these issues. For example, SoT (Ning et al., 2024) and PASTA (Jin et al., 2025) elicit intermediate sub-questions or steps that can be executed concurrently, optimizing both accuracy and efficiency. In contrast to linear decompositions, ToT (Yao et al., 2023) organizes reasoning into a tree structure, allowing deliberate search over candidate thoughts, while GoT (Besta et al., 2024) and S-DAG (Dong et al., 2025) extend this idea to graph-based reasoning, enabling flexible, reusable intermediate nodes with complex dependency patterns.
Budgeted Edge–Cloud Routing and Online Adaptation.
Edge-cloud collaboration often involves routing mechanisms where simpler models handle easier tasks and offload more complex ones to stronger models, or where tasks are decomposed and distributed across multiple models (Wang et al., 2025; Li et al., 2025; Ding et al., 2024; Shao et al., 2025). However, many systems operate at coarse granularity (e.g., query-level or stage-level) and do not explicitly optimize for tight, multi-resource constraints such as latency and API budgets. Cost-aware routing methods, like LLM cascades using mixture-of-thought representations (Yue et al., 2023) and FrugalGPT (Chen et al., 2023), address budget-quality trade-offs by selectively invoking stronger models at various stages. Related work like SplitReason (Akhauri et al., 2025) also explores fine-grained offloading, learning to delegate specific reasoning steps to stronger models. However, many of these learned routers rely on stationary assumptions or fixed thresholds, which can hinder performance when the workload, network conditions, or pricing fluctuate. In contrast, HybridFlow adapts dynamically to these changes, optimizing edge-cloud allocation in real-time for more robust and efficient reasoning.
Algorithm 1 HybridFlow: Dependency-Aware Scheduling with Budget-Adaptive Routing
Input: Query $Q$; edge model $M_{\mathrm{edge}}$; cloud model $M_{\mathrm{cloud}}$; planner $M_{P}$; budgets $C_{\max}$; router $f_{\theta}$; step size $\eta$; threshold params $(\tau_{0},\gamma)$
@@
Stage 1: Decompose and initialize
$(T,E)\leftarrow\mathrm{Decompose}(Q;M_{P})$
; build DAG $G=(T,E)$
Initialize in-degrees $\deg(\cdot)$ on $G$; answers map $A[\cdot]\leftarrow\emptyset$
Initialize ready queue $\mathcal{F}\leftarrow\{t\in T:\deg(t)=0\}$
Initialize cumulative cost $C_{\mathrm{used}}\leftarrow 0$ and dual variable $\lambda\leftarrow 0$
Stage 2: Schedule, route, and execute
while
$\mathcal{F}$ not empty do
   Pop a ready subtask $t$ from $\mathcal{F}$
$x\leftarrow\big(t,\{A[p]\}_{p\in\mathrm{Dep}(t)}\big)$
$\hat{u}\leftarrow\sigma(f_{\theta}(\mathrm{Embed}(t)))$ ;
[…]
 // Eq. ($0,1)\tau\leftarrow\mathrm{clip}(\tau_{0}+\gamma\lambda,$0,1) ;
 // Eq. (11)if $\bar{u}>\tau$ and $C_{\mathrm{used}}+c(t)\leq C_{\max}$ then
    $A[t]\leftarrow M_{\mathrm{cloud}}(x)$;  $C_{\mathrm{used}}\leftarrow C_{\mathrm{used}}+c(t)$
@@
    $A[t]\leftarrow M_{\mathrm{edge}}(x)$
 foreach $s\in\mathrm{Succ}(t)$ do
    $\deg(s)\leftarrow\deg(s)-1$; if $\deg(s)=0$ then
        push $s$ into $\mathcal{F}$
Stage 3: Aggregate
$R\leftarrow\mathrm{Aggregate}(\{A[t]\}_{t\in T}\ \text{in topological order})$
return
$R$
@@ 3 HybridFlow Framework
HybridFlow enables efficient edge–cloud collaborative inference by combining dependency-aware task planning with cost-aware routing. As illustrated in Fig. 1, the system has two tightly connected components. First, an edge-side planner decomposes a complex query into a set of interdependent subtasks and constructs a task-level DAG, enabling each subtask to execute as soon as its prerequisites are resolved. This exposes fine-grained parallelism and reduces wall-clock latency compared with strictly sequential reasoning. Second, a resource-aware router assigns each subtask to either the edge model $M_{\mathrm{edge}}$ or the cloud LLM $M_{\mathrm{cloud}}$ based on its predicted utility and the system’s real-time budget status. By coupling decomposition with budget-constrained routing, HybridFlow explicitly balances accuracy and efficiency. We summarize the notations in Table 4 and outline the full procedure in Algorithm 1.
@@ 4.3 Ablation Study
We evaluate our adaptive router against deterministic and naive allocation strategies (Table 3). Using only the edge model (Edge) incurs zero API cost but yields the lowest accuracy (25.54%), whereas executing all subtasks in the cloud (Cloud) achieves the highest accuracy (57.28%) with higher latency and API cost ($C_{\text{API}}{=}0.0185$). The Random baseline improves accuracy (46.00%) by occasionally offloading, but lacks a principled mechanism to prioritize high-value subtasks, resulting in lower utility. In contrast, HybridFlow attains 53.33% accuracy at a much lower API cost ($C_{\text{API}}{=}0.0075$) and achieves the highest utility score (0.7940), indicating a more favorable accuracy–cost trade-off under our unified metric. Overall, these results suggest that the learned router better aligns subtask-level benefit with resource expenditure.
@@
To characterize the router’s within-query allocation, we count how many subtasks are executed on the edge vs. the cloud at each subtask position. Figure 4 shows a clear position-dependent pattern rather than uniform offloading. At early positions, the adaptive threshold is relatively low and the remaining budget is large, making cloud execution more likely. As reasoning proceeds, the threshold increases and eventually saturates; correspondingly, routing shifts toward the edge, with few cloud calls at later positions. Meanwhile, the total number of subtasks decreases at deeper positions, suggesting that many queries resolve key reasoning steps early. Overall, HybridFlow concentrates cloud usage on early, high-impact subtasks and relies on the edge model for downstream subtasks as the routing criterion becomes stricter.
@@
To isolate the role of the base threshold, we vary the fixed offload threshold $\tau_{0}$ in Eq. 11 without dynamic resource adaptation and measure accuracy, offload rate, latency, and cost. As shown in Figure 4 and Table 5, increasing $\tau_{0}$ makes routing more conservative: the offload rate decreases monotonically from 100.00% ($\tau_{0}{=}0$) to 0.00% ($\tau_{0}{=}1$), and the normalized cost decreases accordingly from c=0.7760c{=}0.7760 to c=0.2000c{=}0.2000 at $\tau_{0}{=}0.9$ (and is undefined when no cloud calls are made). Accuracy declines smoothly over the same range, from 57.28% at $\tau_{0}{=}0$ to 25.54% at $\tau_{0}{=}1$. This consistent trade-off suggests that the predicted utilities $\hat{u}_{i}\in[0,1]$ yield coherent decision boundaries when compared against the normalized cost scale $c_{i}\in[0,1]$.
Notably, the utility score peaks at $\tau_{0}{=}0.6$ with u=0.6329u{=}0.6329, where the router maintains 47.85% accuracy at a moderate normalized cost (c=0.3525c{=}0.3525) and an offload rate of 33.51%. Moving to a looser threshold (e.g., $\tau_{0}{=}0.5$) increases both accuracy (51.62%) and cost (c=0.4145c{=}0.4145) but slightly reduces utility (0.6292), while more aggressive offloading further raises cost and eventually decreases utility (e.g., u=0.5652u{=}0.5652 at $\tau_{0}{=}0.4$ and u=0.4090u{=}0.4090 at $\tau_{0}{=}0$). These results indicate that a single global threshold has a narrow optimal operating region.
In contrast, HybridFlow achieves a substantially higher utility (0.7940 in Table 3) by dynamically adapting routing decisions to the evolving budget, outperforming any fixed-threshold setting while avoiding both excessive offloading and premature underutilization.
@@ Appendix A Notations
$\tau_{t}$ Adaptive routing threshold at time $t$ [Eq. (11)].
@@ B.3 Primal–Dual Dynamics and Adaptive Thresholding
HybridFlow’s adaptive threshold in Eq. (11) of the main text,